Cullumber Corporation issues $550,000 of 9% bonds, due in 9 years, with interest payable semiannually. At the time of issue, the market rate for such bonds is 10%.

Compute the issue price of the bonds. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971.)

Issue price of the bonds $
Answer #1
Issue price of the bonds is $ 5,17,854
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.05)^-18)/0.05 i = 10%/2 = 5.00%
=          11.68959 n = 9*2 = 18
Present value of 1 = (1+i)^-n
= (1+0.05)^-18
=             0.41552
Present value of coupon = $       5,50,000 * 4.50% * 11.68959 = $       2,89,317
Present value of par value = $       5,50,000 * 0.415521 0.41552 = $       2,28,536
Present value of cash flows $       5,17,854